# Early-Exit Graph Neural Networks

### v1

@@ Abstract
Early-exit mechanisms allow deep neural networks to halt inference as soon as classification confidence is high enough, adaptively trading depth for confidence, and thereby cutting latency and energy on easy inputs while retaining full-depth accuracy for harder ones. Similarly, adding early exit mechanisms to Graph Neural Networks (GNNs), the go-to models for graph-structured data, allows for dynamic trading depth for confidence on simple graphs while maintaining full-depth accuracy on harder and more complex graphs to capture intricate relationships. Although early exits have proven effective across various deep learning domains, their potential within GNNs in scenarios that require deep architectures while resisting over-smoothing and over-squashing remains largely unexplored. We unlock that potential by first introducing Symmetric-Anti-Symmetric Graph Neural Networks (SAS-GNN), whose symmetry-based inductive biases mitigate these issues and yield stable intermediate representations that can be useful to allow early exiting in GNNs. Building on this backbone, we present Early-Exit Graph Neural Networks (EEGNNs), which append confidence-aware exit heads that allow on-the-fly termination of propagation based on each node or the entire graph. Experiments show that EEGNNs preserve robust performance as depth grows and deliver competitive accuracy on heterophilic and long-range benchmarks, matching attention-based and asynchronous message-passing models while substantially reducing computation and latency. We plan to release the code to reproduce our experiments.
Department of Computer Science, Control and Management Engineering,
Sapienza University of Rome, Rome, Italy
Institute of Information Science and Technologies "Alessandro Faedo" - ISTI-CNR, Pisa, Italy
Corresponding author: difrancesco@diag.uniroma1.it
§Information Engineering Department, University of Pisa, Pisa, Italy
@@ 1 Introduction
Deep learning models are increasingly deployed in latency and energy-constrained settings (e.g., mobile AR, autonomous drones, real-time recommendation). Graph Neural Networks (GNNs) inherit these constraints because their message-passing depth directly translates into runtime and energy costs. In such scenarios, adapting computational effort to input difficulty is critical for both efficiency and sustainability. Graph Neural Networks (GNNs) have emerged as a powerful class of deep learning models designed to process graph-structured data. Graphs are a natural way to represent information across various domains, such as text, audio, images, knowledge representation, and social networks (Ul Qumar et al., 2023; van den Berg et al., 2017; Hamaguchi et al., 2018). The ability of GNNs to model these complex structures has allowed them to excel in tasks such as node and graph classification (Xu et al., 2019), as well as edge prediction (Kumar et al., 2020). Most GNNs follow the message-passing paradigm (Gilmer et al., 2017), and they are referred to as Message-Passing Neural Networks (MPNNs). In this paradigm, each layer updates the representation of a node by aggregating feature information from its immediate neighbors through graph convolutions, attention mechanisms, or learned neural functions (Veličković, 2023). Consequently, the number of layers of an MPNN dictates the length of paths from which information from nodes can propagate. Intuitively, increasing the number of layers should allow for better integration of long-range information, enabling messages to traverse farther across the graph and enriching each node’s representation with broader structural context. In practice, however, depth is a double-edged sword: too many layers trigger over-smoothing, in which node embeddings become indistinguishable, and over-squashing, where information from distant nodes is aggressively compressed (NT and Maehara, 2019; Topping et al., 2022); too few layers lead to under-reaching, where messages cannot cover the task-specific problem radius (Alon and Yahav, 2021). Because the problem radius cannot be known a priori, the layer count becomes a delicate hyperparameter that should be set to the shallowest depth still able to span the necessary receptive field while keeping model size and training cost as low as possible. However, even if we were able to set the number of layers to its optimal value, it is known that Message-Passing Neural Networks remain no more expressive than the 1-Weisfeiler–Lehman test (Weisfeiler and Lehman, 1968; Xu et al., 2019), limiting their ability to distinguish certain graph structures regardless of depth.
Our idea. Rather than selecting a single “best” depth, we let the network decide on-the-fly: we endow GNNs with early exits so that each node (or the whole graph) can halt message passing as soon as its prediction is confident. We let each node (or the whole graph) decide how far messages need to travel by attaching confidence-aware exit heads to a depth-stable backbone. However, early exits are reliable only if intermediate representations remain both informative and parameter-efficient; achieving this requires a new backbone. To that end, we design Symmetric–Anti-Symmetric GNNs (SAS-GNNs), whose weight-shared, ODE-inspired message passing produces stable embeddings and constant memory usage—making early exit practical at scale.
@@
    EEGNN. First end-to-end early-exit GNN: each node or the whole graph halts message passing when confident via Gumbel–Softmax heads, reducing inference time and removing depth tuning.
@@
    SAS-GNN backbone. A weight-shared, symmetry/anti-symmetry MPNN (Neural ODE-inspired) that offers stable mid-layer states for safe early exits with constant memory.
@@
    Theory. We prove SAS-GNN preserves node-wise information and creates controllable attraction/repulsion, countering over-smoothing and over-squashing—thus supporting reliable exits.
@@
    Results. On heterophilic benchmarks and LRGB, EEGNN/SAS-GNN match or beat attention- and asyncronous-based MPNNs with far fewer parameters, no normalization/dropout, and lower latency.
@@ 2 Related Work
Early-Exit for GNNs. Our approach shares similarities with Spinelli et al. (2021), which introduces an early-exit mechanism for GNNs, allowing nodes to halt their updates during message passing. While their method considers over-smoothing (OST), no solution for over-squashing (OSQ) was proposed. Additionally, their design requires auxiliary loss terms to enable differentiable node-level exit decisions. Other works have explored early exit mechanisms in GNNs, but these do not investigate theoretically either OST or OSQ (Xiao et al., 2021; Han et al., 2024).
In contrast, we are the first to propose early-exit not to bypass OST and OSQ. For instance, our method enables nodes to remain active in the network as long as necessary for feature extraction, naturally addressing both OST and OSQ through its architecture.
Then, our early-exit mechanism is fully differentiable and trained end-to-end using only the task loss, without the need for additional regularization or supervision signals.
While all of the above works considering early-exit have focused solely on node-level tasks, we are the first to include an extension for graph classification/regression.
Asynchronous / Transformer GNNs. Graph Transformers (GTs) (Shi et al., 2021) use self-attention to enable unrestricted node interactions, avoiding reliance on graph topology or increased depth to model long-range dependencies. However, this comes at a cost of quadratic time complexity in the number of nodes, limiting scalability. While GTs seem promising, Cai et al. (2023) shows they are not inherently more expressive than MPNNs with a virtual node, which also allows global communication. Then, in practice, GTs do not consistently outperform classical MPNNs like GCN (Kipf and Welling, 2017) on the LRGB benchmark (Tönshoff et al., 2023), casting doubt on their effectiveness for long-range tasks. Moreover, both GTs and MPNNs often rely on normalization and dropout (Luo et al., 2024), which can obscure the model’s internal behavior. Asynchronous MPNNs (Finkelshtein et al., 2024; Errica et al., 2023) dynamically adapt the graph topology at each layer to track specific paths between distant nodes and improve long-range communication, though at the cost of increased architectural complexity. They are opposed to synchronous MPNNs that operate on a fixed topology. Co-GNN (Finkelshtein et al., 2024) introduces node-level interaction modes (e.g., isolate, broadcast, listen) to address OST and OSQ. Inspired by this, we also allow discrete node decisions—but in our model, we choose whether to exit or continue, preserving the original topology and ensuring efficiency. Unlike Co-GNNs, which require manual depth tuning and are costlier at inference, our method is lightweight and scalable. AMP (Errica et al., 2023) learns depth and a message filtering strategy to address OST, OSQ, and under-reaching, but applies a fixed depth at inference. In contrast, our model adapts depth at both training and test time while using a fixed topology.
Neural ODEs methods. A smaller subset of MPNNs belongs to the class of Graph Neural Ordinary Differential Equations (ODEs) (Poli et al., 2021). These approaches aim to design more principled architectures by avoiding components like dropout and normalization layers, whether not supported by theory. They also tend to be memory-efficient, as weight matrices are typically shared across layers. Although models in this family have been proposed to address OST or OSQ individually (Di Giovanni et al., 2023; Gravina et al., 2023), to the best of our knowledge, no method has tackled both phenomena simultaneously within this framework, neither have they been used to design early-exit GNN architectures.
An extended version of the related works, covering GNNs in general and Early-Exit Neural Networks, is in Appendix G.
@@ 3 Methodology
We use bold fonts for both matrices and vectors, with uppercase letters representing matrices and lowercase letters representing vectors (e.g., $\mathbf{M}$, $\mathbf{v}$). Scalars are denoted by italic letters (e.g., s). Notation. Let $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathbf{X})$ be an undirected graph, where $\mathcal{V}$ is the set of nodes, $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ is the set of edges and $\mathbf{X}\in\mathbb{R}^{n\times m}$ is the instance matrix containing vector representations of features for each node. The $u$-$th$ row of the instance matrix is represented by $\mathbf{x}_{u}\in\mathbb{R}^{m}$. The number of nodes in $\mathcal{G}$ is denoted by $|\mathcal{V}|=n$, and the number of edges by $|\mathcal{E}|$. The symbol $\Gamma(u)$ represents the neighborhood of node $u$, and $|\Gamma(u)|$ denotes its degree. The diagonal matrix $\mathbf{D}\in\mathbb{R}^{n\times n}$ is such that the $u$-$th$ element on the diagonal is equal to $|\Gamma(u)|$. The set of edges $\mathcal{E}$ can also be expressed as the adjacency matrix $\mathbf{A}\in\{0,1\}^{n\times n}$, where $A_{uv}=1$ if $(u,v)\in\mathcal{E}$, and $A_{uv}=0$ otherwise. Graph Neural Networks. Most GNNs utilize the message-passing formalism to extract features from graph-structured datasets (Gilmer et al., 2017). In this formalism, the instance matrix $\mathbf{X}$ is iteratively transformed across layers during the forward pass of the GNN. We define the initial hidden representation as $\mathbf{H}^{0}=\mathbf{X}$, or as $\mathbf{H}^{0}=f(\mathbf{X})$ when using a learnable projection function $f$, such as a multi-layer perceptron (MLP). At each layer $l$, the model computes an intermediate representation matrix $\mathbf{H}^{l}$ for $0\leq l\leq L$, where the feature vector of node $u$ is denoted by $\mathbf{h}_{u}^{l}\in\mathbb{R}^{m^{\prime}}$, with $m^{\prime}$ representing the hidden dimension. This process is conditioned at each layer by $\mathbf{A}$ up to layer $L$, yielding the final node representations $\mathbf{H}^{L}\in\mathbb{R}^{n\times m^{\prime}}$, which we denote as $\mathbf{Z}$. Once we have $\mathbf{Z}$, it is used for downstream tasks such as node or graph classification, typically through a specific readout function, followed by an MLP $g(\cdot)$. The former is given by $\hat{y}_{v}=g(\mathbf{z}_{v})$ and the latter by $\hat{y}=g(\text{Pool}(\mathbf{Z}))$, where Pool (e.g.,  mean or max) is any permutation-invariant aggregator. When considering multiple graphs, we denote each graph $i$ in a dataset $\mathcal{D}=\{\mathcal{G}_{i}:i\in\mathcal{D}\}$ where $\mathcal{G}_{i}=(\mathcal{V}_{i},\mathcal{E}_{i},\mathbf{X}_{i})$, with associated adjacency matrix $\mathbf{A}_{i}$. Similarly, we define the hidden representation matrix at layer $l$ for graph $i$ as $\mathbf{H}^{l}_{i}$. When possible, we describe the approach for a single graph. We define a task as transductive when it is performed on a single graph, while the task is inductive when datasets contain more than one graph, namely those in the LRGB.
@@
Our approach. A naïve scheme would let every node draw an $\arg\max$ over the two actions—exit or continue—at each layer and stop if it chooses the first. Unfortunately, the hard $\arg\max$ is non-differentiable, so the exit policy cannot be learned with standard back-propagation.
To retain end-to-end differentiability, we therefore need a soft, trainable substitute for the discrete decision. This requirement leads to two concrete design goals:
@@
Because early-exit heads can only act on reliable hidden states, O1 is a prerequisite for O2.
Symmetric-Anti-Symmetric Graph Neural Network (O1).
To accomplish O1, we build upon two message-passing schemes proposed in recent years. One is the Anti-Symmetric Deep Graph Network (A-DGN) (Gravina et al., 2023), and the other is the Gradient Flow Framework (GRAFF) (Di Giovanni et al., 2023); full details are provided in Appendix B. The former was introduced to prevent OSQ by leveraging antisymmetric learnable matrices. This results in a GNN that behaves as a stable and non-dissipative dynamical system—i.e., the real parts of each Jacobian’s eigenvalue are zero, namely non-positive (stability), and non-negative (non-dissipativeness). Concretely, this ensures that the gradient $\partial\mathbf{h}_{i}^{l}/\partial\mathbf{h}_{i}^{0}$ for each node $i$ remains well-conditioned across layers, neither vanishing nor exploding. This enables long-range dependencies to be preserved during propagation, which is essential for tasks where the problem radius is large. Further discussion is available in Appendix B. The latter, instead, was introduced to deal with node classification in heterophilic graphs [^1] A graph is heterophilic when adjacent nodes tend to share different class labels.. Since OST causes nodes to converge to the same representation, its negative effect is more enhanced when learning in heterophilic graphs, where adjacent nodes with different classes should have a distinguished representation. GRAFF takes advantage of symmetric learnable matrices, which provably enable attraction and repulsion edge-wise to prevent adjacent nodes from becoming similar in the limit of many layers. Since these approaches rely uniquely on the parameter design level, our theoretical contribution is to devise a unique architecture that takes advantage of both methods. Specifically, we designed a message-passing scheme that accounts for OST through symmetric weight matrices and OSQ through the antisymmetric counterpart.
We name such architecture Symmetric-Anti-Symmetric Graph Neural Network (SAS-GNN). Its message-passing rule is
@@
where we have, without loss of generalization, $\mathbf{H}^{t}\equiv\mathbf{H}^{l}$, $\mathbf{\bar{A}}=\mathbf{D}^{-\frac{1}{2}}\mathbf{A}\mathbf{D}^{-\frac{1}{2}}$, that is the normalized adjacency matrix. We refer to $\sigma_{1},\sigma_{2}$ as the non-linear activation functions, and finally $\mathbf{\Omega}_{as}$, $\mathbf{W}_{s}\in\mathbb{R}^{m^{\prime}\times m^{\prime}}$ are the antisymmetric and symmetric trainable weight matrices. We formalised Equation (1) as an ODE since we build upon the Graph Neural ODE framework. We can integrate it via the Euler discretization as:
@@
Here, $\tau$ is the integration step. Weight matrices are shared across layers, as specified in GRAFF. A-DGN performs similarly with or without weight sharing (Gravina et al., 2023), but we adopt it for space efficiency when scaling to many layers ($t\rightarrow\infty$). This design is supported by the following theorems.
@@
Let us assume that the node features $\mathbf{H}^{t}$ evolve according to Equation (1). Assuming that $\bar{\mathbf{A}}$ does not contain self-loops, and the derivative of $\sigma_{1}$ is bounded, then the evolution of $\mathbf{H}^{t}$ is stable and non-dissipative.
Proof sketch. This result follows from a standard stability analysis of the ODE in Equation (1). The antisymmetric term $-\sigma_{2}(\mathbf{H}^{t}\boldsymbol{\Omega}_{\mathrm{as}})$ contributes Jacobian eigenvalues with purely imaginary components, while the symmetric term involving $\mathbf{W}_{s}$ does not increase their real parts.
@@
Let us assume that the node features $\mathbf{H}^{t}$ evolve according to Equation (1). Assuming that $\sigma_{1}$, and $\sigma_{2}$ are defined s.t. $\forall x\in\mathbb{R}$, $\sigma_{1}(x),\sigma_{2}(x)\geq 0$. The evolution of $\mathbf{H}^{t}$ minimizes a parameterized energy functional $E_{\theta}(\mathbf{H}^{t})$, inducing attraction or repulsion among adjacent nodes.
The functional that we refer to is the following
Proof sketch. We show that Equation (1) monotonically decreases the energy functional in Equation (3). The symmetry of $\mathbf{W}_{s}$ allows its decomposition into eigencomponents, revealing how the dynamics induce attraction or repulsion between nodes. The edge-wise interpretation is detailed in Equation 28 (Appendix).
@@
Proof sketch. The proof, follows directly from Theorems 3.1 and 3.2, as ReLU+TanH and ReLU satisfy the required boundedness and non-negativity conditions. We adopt this activation pair in all experiments, and show in Appendices F.3 and F.6 that it impacts the performance.
Since the LRGB made available edge features for each graph, we also propose a SAS-GNN version that encompasses their use $\mathbf{E}\in\mathbb{R}^{|\mathcal{E}|\times d}$ as follows.
@@
Let us assume that the node features $\mathbf{H}^{t}$ evolve according to equation (4). If $\sigma_{1}(x)=\text{ReLU}(\text{tanh}(x))$, $\sigma_{2}(x)=\text{ReLU}(x)$, and $f_{e}(\mathbf{E})\equiv-ReLU(\mathbf{B}\mathbf{E}\mathbf{W}_{e})$, then the evolution of $\mathbf{H}^{t}$ is stable and non-dissipative and minimizes a parameterized energy functional $E_{\theta}(\mathbf{H}^{t})$, inducing attraction or repulsion among adjacent nodes.
This theorem can be proved equivalently to the previous ones. Theorems’ proofs are included in Appendix C.
Algorithm 1 Neural Adaptive-Step Early-Exit GNNs for Node Classification
1:Initialize $\mathbf{H}^{0}$, $L$, $\mathbf{\bar{A}}$, $f_{e}$, $f_{c}$, $f_{\nu}$, $\mathbf{W}_{s}$, $\mathbf{\Omega}_{as}$, $\sigma_{1}$, $\sigma_{2}$, $\nu_{0}$, $\mathbf{Z}=\mathbf{0}_{n\times d}$, $exit\_list=\{\}$
2:for $l=0$ to $L$ do
3:     $\mathbf{C}^{l}\leftarrow f_{c}(\mathbf{H}^{l})$; $\bm{\nu}^{l}\leftarrow f_{\nu}(\mathbf{H}^{l},\nu_{0})$
4:     $\mathbf{c}^{l}\leftarrow gumbel\_softmax(\mathbf{C}^{l},\bm{\nu}^{l})$
5:     $\bm{\tau}^{l}\leftarrow\mathbf{c}^{l}(\texttt{0})$
6:     $\mathbf{H}^{l+1}\leftarrow\mathbf{H}^{l}+\bm{\tau}^{l}\sigma_{1}(-\sigma_{2}(\mathbf{H}^{l}\mathbf{\Omega}_{as})+f_{e}(\mathbf{E})+\mathbf{\bar{A}}\mathbf{H}^{l}\mathbf{W}_{s})$
7:     for $i=0$ to $n$ do
8:         if $argmax\{\mathbf{c}^{l}\}=\texttt{1}\wedge i\notin exit\_list$ then
9:              $\mathbf{Z}_{i}\leftarrow\mathbf{h}_{i}^{l}$; $exit\_list.add(i)$
10:for $i=0$ to $n$ do
11:     if $i\notin exit\_list$ then
12:         $\mathbf{Z}_{i}\leftarrow\mathbf{h}_{i}^{L}$;$exit\_list.add(i)$
13:return $\mathbf{Z}$
Gumbel Softmax Early-Exit Mechanism (O2).

### v2

@@ Abstract
Early-exit mechanisms allow deep neural networks to stop inference once prediction confidence is high, reducing latency and energy on easy inputs while retaining full-depth accuracy on harder ones. Similarly, adding early exit mechanisms to Graph Neural Networks (GNNs), the go-to models for graph-structured data, allows for dynamic trading depth for confidence on simple graphs while maintaining full-depth accuracy on harder ones to capture intricate relationships. Yet, their potential in deep GNNs, where over-smoothing, over-squashing or more generally vanishing gradients prevent these model to properly learn, remains largely unexplored. To address this, we introduce Symmetric-Anti-Symmetric GNNs (SAS-GNN), whose symmetry-based inductive biases yield stable intermediate representations that support safe early exits. Building on this backbone, we propose Early-Exit GNNs (EEGNNs), which attach confidence-aware exit neural heads which are trainable end-to-end based on the task objective, enabling on-the-fly termination at node or graph level. Experiments show that EEGNNs learn task-driven exit strategies, while achieving competitive results on heterophilic graphs and long-range tasks. Even when not outperforming the strongest baselines, EEGNNs consistently deliver favorable accuracy–efficiency trade-offs thanks to their adaptive and parameter-efficient design. We plan to release the code to reproduce our experiments.
Machine Learning
@@ 1 Introduction
Deep learning models are increasingly deployed in latency and energy-constrained settings (e.g., mobile AR, autonomous drones, real-time recommendation). Graph Neural Networks (GNNs) inherit these constraints because their message-passing depth directly translates into runtime and energy costs. In such scenarios, adapting computational effort to input difficulty is critical for both efficiency and sustainability. GNNs process graph-structured data across domains such as text, images, knowledge graphs, and social networks (Ul Qumar et al., 2023; van den Berg et al., 2017; Hamaguchi et al., 2018), excelling in tasks like node/graph classification (Xu et al., 2019) and link prediction (Kumar et al., 2020). Most GNNs are Message-Passing Neural Networks (MPNNs) (Gilmer et al., 2017), where each layer updates node states by aggregating neighbor features via convolutions, attention, or learned functions (Veličković, 2023). The number of layers thus determines how far information can propagate: deeper models capture broader context but incur higher cost. Intuitively, increasing the number of layers should allow for better integration of long-range information, enabling messages to traverse farther across the graph and enriching each node’s representation with broader structural context. In practice, however, depth is a double-edged sword: on the one hand, too many layers cause over-smoothing (NT and Maehara, 2019), where node embeddings become indistinguishable, and over-squashing, where information from distant nodes is severely compressed due to topological or computational bottlenecks (Alon and Yahav, 2021; Topping et al., 2022; Arnaiz-Rodriguez and Errica, 2025), both hindering the success of MPNNs on tasks requiring long-range propagation; on the other hand, too few layers result in under-reaching, where messages fail to cover the task-specific problem radius (Alon and Yahav, 2021). Because the problem radius cannot be known a priori, the layer count becomes a delicate hyperparameter that should be set to the shallowest depth still able to span the necessary receptive field while keeping model size and training cost as low as possible.
Our idea. Rather than fixing a single “best” depth, we allow GNNs to decide on-the-fly: each node or the entire graph can halt message passing once its prediction is sufficiently confident. However, as shown in Figure 1, classic MPNNs do not reliably scale in depth due to the above issues, even in the presence of residual connections (Scholkemper et al., 2025). As a consequence, early-exit mechanisms cannot be straightforwardly applied to regimes that inherently require deep GNNs, such as long-range reasoning, unless we propose specialized approaches. When depth itself is unreliable, early exits merely act as a safeguard, dedicating time to easy inputs, but ignoring harder samples, which contrasts with the core purpose of early-exit (Scardapane et al., 2020). This limitation was already evident in the seminal work of Spinelli et al. (2021), where the exit mechanism was learned via maximum budget loss function. Reliable early exits require intermediate representations that remain informative across layers, so that the model can safely continue message passing when it is not confident yet. However, this requirement is not met by current methodologies: the trade-off between exiting early and processing features deeply enough is fundamentally unstable. In fact, we show that existing early-exit, when rely on budget-aware training objectives bias the model toward premature halting focusing more on computational constraints rather than predictive correctness. In contrast to this, we believe that a reliable early-exit mechanism must be agnostic to any predefined budget, and allow for exit coherently with the task objective. To achieve this, we design Symmetric–Anti-Symmetric GNNs (SAS-GNNs), whose weight-shared, ODE-inspired message passing yields informative embeddings and constant memory necessary to deploy EEGNN, an early exit model which learns to halt propagation based on the task.
@@
    EEGNN. The first end-to-end early-exit GNN: nodes or graphs halt when confident via Gumbel–Softmax heads enabling end-to-end training and removing the necessity for depth or budget tuning.
@@
    SAS-GNN backbone. A weight-shared, symmetry/anti-symmetry MPNN (ODE-inspired) that provides stable intermediate states for safe early exits with constant memory.
@@
    Theory. We prove SAS-GNN preserves node information while inducing adaptive attraction/repulsion edge-wise, resulting as a good proxy for deep feature processing and supporting long-range tasks.
@@
    Results. Extensive experiments on heterophilic and long-range benchmarks demonstrate that EEGNN and SAS-GNN match or surpass complex Attention-based and Asynchronous MPNNs. Notably, they achieve this with significantly fewer parameters, no normalization or dropout, and contained latency, while successfully adapting exit strategies to the specific demands of each task.
Figure 1: Classic MPNNs + residual design is not enough for going deep. Texas Dataset.
@@ 2 Related Work
Early-Exit for GNNs. Our approach shares similarities with Spinelli et al. (2021), which introduces the idea of adaptive propagation node-wise for GNNs, allowing nodes to halt their updates during message passing.
While their method discuss depth-related issues like over-smoothing (OST), no solution for designing deep architectures and dealing with long-range tasks has been proposed. Additionally, their design requires auxiliary loss terms to enable differentiable node-level exit decisions.
Other early-exit methods for GNNs (Xiao et al., 2021; Han et al., 2024; Abbahaddou et al., 2025) are restricted to node-level tasks and do not investigate deep regimes or OST/OSQ. In contrast, we propose the first fully differentiable mechanism that naturally addresses deep MPNN flaws without auxiliary supervision. Furthermore, we are the first to extend this capability to graph classification and regression tasks.
Asynchronous / Transformer GNNs. Graph Transformers (Shi et al., 2021) capture long-range dependencies via global attention but incur quadratic computational costs. Asynchronous MPNNs (Finkelshtein et al., 2024; Errica et al., 2023) attempt to reduce this cost by adapting topology at each layer at the cost of an increased architectural complexity w.r.t. classic MPNNs. Unlike these approaches, our model adapts depth dynamically at both train and test time without altering the graph topology, sharing the complexity with classic MPNNs.
Neural ODE methods. Graph Neural ODEs (GraphNODEs) (Poli et al., 2021) model feature propagation as a continuous-time dynamical system, providing a rigorous framework to analyze stability. While initially used to address over-smoothing (OST) (Rusch et al., 2022; Di Giovanni et al., 2023), recent works leverage physics-inspired architectures to mitigate over-squashing (OSQ) and enhance long-range propagation (Gravina et al., 2023, 2025; Trenta et al., 2025).
In this work, we draw inspiration from prior approaches addressing both OST and OSQ, and we propose EEGNN, the first GraphNODE who natively supports an early-exit mechanism. We provide an extended related work discussion in Appendix G.
@@ 3 Methodology
We use bold fonts for both matrices and vectors, with uppercase letters representing matrices and lowercase letters representing vectors (e.g., $\mathbf{M}$, $\mathbf{v}$). Scalars are denoted by italic letters (e.g., s).
Notation. Let $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathbf{X})$ be an undirected graph, where $\mathcal{V}$ is the set of nodes, $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ is the set of edges and $\mathbf{X}\in\mathbb{R}^{n\times m}$ is the instance matrix containing node features representations. The $u$-thth row of the instance matrix is represented by $\mathbf{x}_{u}\in\mathbb{R}^{m}$. $\mathcal{G}$ has $n=|\mathcal{V}|$ nodes and $|\mathcal{E}|$ edges. $\Gamma(u)$ represents the neighborhood of node $u$, and $|\Gamma(u)|$ denotes its degree. The degree matrix $\mathbf{D}\in\mathbb{R}^{n\times n}$ has diagonal entries $D_{uu}=|\Gamma(u)|$. The set of edges $\mathcal{E}$ can also be expressed as the adjacency matrix $\mathbf{A}\in\{0,1\}^{n\times n}$, where $A_{uv}=1$ if $(u,v)\in\mathcal{E}$, and $A_{uv}=0$ otherwise.
Graph Neural Networks. Most GNNs follow the message-passing paradigm (Gilmer et al., 2017), where node features are iteratively updated. Let $\mathbf{H}^{0}=\mathbf{X}$, or $\mathbf{H}^{0}=f(\mathbf{X})$ when using a learnable projection (e.g., an MLP). At layer $l$, the hidden matrix is $\mathbf{H}^{l}$ with node embeddings $\mathbf{h}_{u}^{l}\in\mathbb{R}^{m^{\prime}}$, and the process continues up to $L$, conditioned on $\mathbf{A}$. The final representations $\mathbf{H}^{L}=\mathbf{Z}\in\mathbb{R}^{n\times m^{\prime}}$ are used for downstream tasks. For node classification, $\hat{y}_{v}=g(\mathbf{z}_{v})$, and for graph classification, $\hat{y}=g(\text{Pool}(\mathbf{Z}))$, where Pool is a permutation-invariant operator (e.g., mean, max). For datasets $\mathcal{D}={\mathcal{G}_{i}}$ with graphs $\mathcal{G}_{i}=(\mathcal{V}_{i},\mathcal{E}_{i},\mathbf{X}_{i})$ and adjacency $\mathbf{A}_{i}$, we analogously define $\mathbf{H}^{l}_{i}$. We use single-graph notation when possible. We define a task as transductive when it is performed on a single graph, while the task is inductive when datasets contain more than one graph.
@@
Our approach. A naïve scheme would let every node draw an $\arg\max$ over the two actions—exit or continue—at each layer and stop if it chooses the first. Since the hard $\arg\max$ is non-differentiable, the exit policy cannot be learned with standard back-propagation. To retain end-to-end differentiability, we need a soft, trainable substitute for the discrete decision. This requirement leads to two concrete design goals:
@@
Because we want early-exit heads to act on reliable hidden states, O1 is a prerequisite for O2.
Symmetric-Anti-Symmetric Graph Neural Network (O1). To accomplish O1, we build upon two message-passing schemes proposed in recent years. One is the Anti-Symmetric Deep Graph Network (A-DGN) (Gravina et al., 2023), and the other is the Gradient Flow Framework (GRAFF) (Di Giovanni et al., 2023). A-DGN preserves long-range information by using antisymmetric learnable matrices, yielding a stable and non-dissipative dynamical system where gradients remain well-conditioned across layers, resulting into a good proxy for OSQ. GRAFF, instead, was introduced to deal with node classification in heterophilic graphs [^1] A graph is heterophilic when adjacent nodes tend to share different class labels.. Since OST causes nodes to converge to the same representation, its negative effect is enhanced when learning in heterophilic graphs. GRAFF takes advantage of symmetric learnable matrices, that induce attraction and repulsion edge-wise to prevent adjacent nodes from becoming similar in the limit of many layers. Further discussions on GRAFF and A-DGN are available in Appendix B.
Building on these ideas, we design the Symmetric–Anti-Symmetric GNN (SAS-GNN), where antisymmetric matrices preserve long-range dependencies and symmetric ones enforce discriminability across classes. Its message-passing rule is
@@
$\mathbf{\bar{A}}=\mathbf{D}^{-\frac{1}{2}}\mathbf{A}\mathbf{D}^{-\frac{1}{2}}$ is the normalized adjacency matrix, $\sigma_{1},\sigma_{2}$ are non-linear activation functions, and finally $\mathbf{\Omega}_{as}$, $\mathbf{W}_{s}\in\mathbb{R}^{m^{\prime}\times m^{\prime}}$ are the antisymmetric and symmetric trainable weight matrices. We formalized Equation (1) as an ODE since we build upon the Graph Neural ODE framework. We integrate it via the Euler discretization as:
@@
Here, $\tau$ is the integration step. Weight matrices are shared across layers, as specified in GRAFF. A-DGN performs similarly with or without weight sharing (Gravina et al., 2023), but we adopt it for space efficiency when scaling to many layers ($t\rightarrow\infty$).
This design is supported by the following theorems.
@@
Let us assume that the node features $\mathbf{H}^{t}$ evolve according to Equation (1). if $\bar{\mathbf{A}}$ does not contain self-loops, and the derivative of $\sigma_{1}$ is bounded, then the evolution of $\mathbf{H}^{t}$ is stable and non-dissipative.
Proof sketch. This result follows from a standard stability analysis of the ODE in Equation (1). The antisymmetric term $\sigma_{2}(\mathbf{H}^{t}\boldsymbol{\Omega}_{\mathrm{as}})$ contributes Jacobian eigenvalues with purely imaginary components, while the symmetric term involving $\mathbf{W}_{s}$ does not increase their real parts.
We can interpret the dynamics of SAS-GNN through an energy functional
The functional in Eq. (3) naturally encodes both attractive and repulsive forces between adjacent nodes through the symmetric weights $W_{s}$. Minimization therefore drives representations to respect structural patterns such as homophily (attraction) or heterophily (repulsion). More formally the following theorem holds.
@@
Let us assume that the node features $\mathbf{H}^{t}$ evolve according to Equation (1). Assuming that $\sigma_{1}$, and $\sigma_{2}$ are defined s.t. $\forall x\in\mathbb{R}$, $\sigma_{1}(x),\sigma_{2}(x)\geq 0$. The evolution of $\mathbf{H}^{t}$ minimizes a parameterized functional $E_{\theta}(\mathbf{H}^{t})$, inducing attraction or repulsion among adjacent nodes.
The proof can be found in Appendix C.
@@
Proof sketch. The proof, follows directly from Theorems 3.1 and 3.2, as ReLU+TanH and ReLU satisfy the required boundedness and non-negativity conditions.
We adopt this activation pair in all experiments, and show in Appendices F.3 and F.6 that it affects the model performances differently w.r.t. other common activation functions.
Since edge features can also be included to improve graph learning (Hu et al., 2020), we also propose a SAS-GNN version that encompasses their use $\mathbf{E}\in\mathbb{R}^{|\mathcal{E}|\times d}$ as follows.
@@
Let us assume that the node features $\mathbf{H}^{t}$ evolve according to equation (4). If $\sigma_{1}(x)=\text{ReLU}(\text{tanh}(x))$, $\sigma_{2}(x)=\text{ReLU}(x)$, $f_{e}(\mathbf{E})\equiv-ReLU(\mathbf{B}\mathbf{E}\mathbf{W}_{e})$, and $\bar{\mathbf{A}}$ does not contain self-loops, then the evolution of $\mathbf{H}^{t}$ is stable and non-dissipative and minimizes a parameterized energy functional $E_{\theta}(\mathbf{H}^{t})$, inducing attraction or repulsion among adjacent nodes.
This theorem can be proved analogously to the previous ones, with full proofs provided in Appendix C. Ablation studies of SAS-GNN with and without edge features are available in Appendix F.5. Empirical validation of SAS-GNN, using metrics for over-smoothing, over-squashing, and performance on long-range tasks and highly heterophilic datasets, is presented in Appendix F.1.
Algorithm 1 Neural Adaptive-Step Early-Exit GNNs
[…]